OpenStudy (anonymous):

Solve the system 5x+7y=16(21y-15x=-48)

OpenStudy (anonymous):

Ummmmm its 5x+7y=16 -21y-15x=-48

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

7y + 5x = 16 -21y - 15x = -48 ------------------ 21y + 15x = 48 -21y - 15x = 48 ============= 0 + 0 = 0

OpenStudy (anonymous):

so what would it be?

OpenStudy (anonymous):

Im confused I thought you suppose to add like terms together

OpenStudy (anonymous):

that's substitution method

OpenStudy (anonymous):

I used elimination method

OpenStudy (anonymous):

well it is too complicated to use the substitution method because if you do there will be too many fractions so I used the elimination method which will eliminate the x or y.

OpenStudy (anonymous):

Oh i really dont know the elimination method

OpenStudy (anonymous):

oh you didn't learn that yet?

OpenStudy (anonymous):

ok if you didn't Ill teach you

OpenStudy (anonymous):

No sorry. My teacher never taught us that

OpenStudy (anonymous):

ok there are 2 types of methods substitution and elimination method

OpenStudy (anonymous):

you already know the substitution method right?

OpenStudy (anonymous):

Sort of

OpenStudy (anonymous):

ok if you are confused do you have any other problem you don't get? so we can use that as an example?

OpenStudy (anonymous):

5x+7y-2z=16 2x+3y-3z=4 -2x-8y-4z=15

OpenStudy (anonymous):

is that one whole question?

OpenStudy (anonymous):

Yeah i have to solve for x,y,and z

OpenStudy (anonymous):

ok lets see we cannot use substitution on this one so lets use elimination method.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well maybe we can

OpenStudy (anonymous):

but will take a very long time

OpenStudy (anonymous):

Well sorry but we cannot use elimination method

OpenStudy (anonymous):

we will use substitution

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

We will find the X first 5x+7y-2z=16 2x+3y-3z=4 -2x-8y-4z=15 ------------------- 5x + 7y - 2z = 16 y = -2/3x + z + 4/3 y = -2/8x + -4/8z + -15/8 --------------------------- so we plug it into the equation of 5x + 7y - 2z = 16 1.) 5x + 7(-2/3x + z + 4/3) = 16 2.) 5x + 7(-2/8x + -4/8z -15/8) = 16 -------------------------------- we are going to multiply The first equation by 3 so we can get whole numbers instead of fraction 1.)5x - 14/3x + 7z + 28/3 = 16 15x - 14x + 21z + 28 = 48 -------------------------- x + 21z + 28 = 48 x + 21z = 20 x = -21z + 20 ------------------- we found the x so we plug it into the x from the beginning of the equation 5x+7y-2z=16 5(-21z + 20) + 7y - 2z -105z + 100 + 7y - 2z = 16 -107z + 7y = -84 7y = -107z - 84 y = -107/7z - 12 ----------------------- plug in the x and the y into the equation 5x+7y-2z=16 5(-21z + 20) + 107z - 12 - 2z = 16 -105z + 100 + 107z - 12 - 2z = 16 z = no solution x = -21z + 20 y = -107/7z - 12

OpenStudy (anonymous):

that was a lot of work

OpenStudy (anonymous):

Yeah thanks for the help on this one. I have another on like this so now i will be able to handle it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

thanks